Cost of Common Equity with and without Flotation

The Evanec Company's next expected dividend, D1, is $3.18; its growth rate is 6%; and its common stock now sells for $36. New stock (external equity) can be sold tonet $32.40 a share.

What is Evanec's percentage flotation cost, F?
Answer #2

a.Cost of retained earnings = (D1/Current price)+Growth rate

 

=($3.18 / $36 )+0.06

= 0.0883+ 0.06

= 0.1483

=14.83%

 

b.% flotation cost= (Current price - Net price)/Current price

 

=(36 – 32.40)/ 36

 

=10%

 

c. Cost of new common stock= (D1/Net price) + Growth rate

 

(3.18/32.40) + 0.06

= 0.0981 + 0.06

= 0.1581

=15.81 %


answered by: Sherly JN-Baptiste
Answer #1
[REMOVED]
Answer #3

The percentage flotation cost (F) can be calculated using the following formula:

F = (P0 - N) / P0

Where: P0 is the current stock price N is the net proceeds per share from the new stock issuance

Given: P0 = $36 N = $32.40

Substituting the values into the formula:

F = ($36 - $32.40) / $36 F = $3.60 / $36 F = 0.1

To express the flotation cost as a percentage, we multiply by 100:

F = 0.1 * 100 F = 10%

Therefore, Evanec's percentage flotation cost (F) is 10%.


answered by: mervetokaz